MILLIMETER AND SUBMILLIMETER-WAVE SPECTRUM OF AMINOACETONITRILE (NH_{2}CH_{2}CN)

\underline{YUTA MOTOKI}, YUKARI TSUNODA, HIROYUKI OZEKI, Department of Environmental Science, Toho University, 2-2-1 Miyama, Funabashi, 274-8510, Japan; KAORI KOBAYASHI, Department of Physics, University of Toyama, 3190 Gofuku, Toyama, 930-8555, Japan.

Aminoacetonitrile (NH_{2}CH_{2}CN) draws an attention with respect to a synthesis of amino acid in interstellar space, as it is considered to be a direct precursor of the simplest amino acid, glycine. Recently the molecule was detected toward SgrB2(N) ^{a}, in which spectral line frequencies are predicted on the basis of the previous millimeter-wave study^{b}. We have measured pure rotational spectrum of aminoacetonitrile in the millimeter and the submillimeter-wave region. About 300 spectral lines including both a-type and b-type transitions were recorded up to 661 GHz, and centrifugal distortion constants up to the octic term were precisely determined. We had to reinvestigate most of the b-type transition lines due to wrong assignments in the previous study. The updated frequency catalogue of the aminoacetonitrile are now accurate enough for astronomically search up to 1000 GHz.

aA. Belloche et al., A&A, 482, 179 (2008)

bM. Bogey, H. Dubus and J. C. Guillmin, J. Mol. Spectrosc., 143, 180 (1990)